srividya get a certain amount of money on retirement.she depoists half the money in the bank gives her daughter half of the remaining and an additional sum of rs 6000.she gives to his the remaining amount which is found to be half of what she gave her daughter.what is the amount srividya received on retirement.?
Answer by advanced_Learner(501)

she recived 72,000 rbs.
let x be the retirement money
she deposited x/2 in the bank
remainder is x-x/(2) which is x/2
she gave 0.5((x)/(2))+6000 to her daughter
remainder is x/2-[0.5((x)/(2))+6000]
remainder is half what the daughter gets
x/2-[0.5((x)/(2))+6000]=0.5(0.5((x)/(2))+6000 )
x=720000
check
720000
half in the bank 36000
remainder is 36000
daughter takes 18000+6000=24000
remainder is 36000-24000=12000 which is half of 24000
cheers
feedback